# Uplift modeling with continuous treatments: A predict-then-optimize approach

### v1

@@ Acknowledgements
This work was supported by Acerta Consult the Flemish Government, through Flanders Innovation & Entrepreneurship (VLAIO, project HBC.2021.0833).
@@ Appendix A Notation overview
Appendix B Details regarding semi-synthetic data
The original IHDP dataset contains 747 observations and 25 features. Following Nie et al. (2021), synthetic counterfactuals are generated by Equations 15-18 where $S_{\mathrm{con}}=\bigl{\{}1,2,3,5,6\bigr{\}}$ is the index set of continuous features, $S_{\mathrm{bin},1}=\bigl{\{}4,7,8,9,10,11,12,13,14,15\bigr{\}}$ and $S_{\mathrm{bin},2}=\bigl{\{}16,17,18,19,20,21,22,23,24,25\bigr{\}}$ are two sets of binary features, $c_{1}=\mathbb{E}\left[\frac{\sum_{i\in S_{\mathrm{dis},1}}x_{i}}{|S_{\mathrm{bin},1}|}\right]$, and $c_{2}=\mathbb{E}\left[\frac{\sum_{i\in S_{\mathrm{dis},2}}x_{i}}{|S_{\mathrm{bin},2}|}\right]$.
@@ Appendix B Details regarding semi-synthetic data
Appendix C Number of dose bins $\delta$
@@
Appendix D Hyperparameter tuning
Method Hyperparameter Values #Models S-Learner (rf) Estimators {20, 50, 100, 200} 48 Criterion {Sq. error, Abs. error} Max depth {5, 15, None} Max features {Sqrt, Log2} S-Learner (mlp) Learning rate {0.0001, 0.001} 32 L2 regularization {0.0, 0.1} Batch size {64, 128} Num layers {2, 3} Hidden size {32, 64} Steps {5000} Optimizer {Adam} DRNet Learning rate {0.0001, 0.001} 64 L2 regularization {0.0, 0.1} Batch size {64, 128} Num repr. layers {2, 3} Num inf. layers {1, 2} Hidden size {32, 64} Steps {5000} Num dose strata {10} Optimizer {Adam} VCNet Learning rate {0.001, 0.01} 8 Batch size {64, 128} Hidden size {32, 64} Steps {5000} Optimizer {Adam}
Table 5: Hyperparameter search space
Appendix E Experiment 1: Dose-response estimation

### v2

@@ Acknowledgements
This work was supported by Acerta Consult and the Flemish Government, through Flanders Innovation & Entrepreneurship (VLAIO, project HBC.2021.0833).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT and Grammarly in order to refine the phrasing of the manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
@@ Appendix A Notation overview
Appendix B Assumptions and mathematical justification of CADE identification
Under the Rubin-Neyman potential outcomes framework, the CADE $\tau_{s}(\mathbf{x})=\mathbb{E}[Y(s)-Y(0)\mid\mathbf{X}=\mathbf{x}]$ quantifies the expected outcome of a continuous treatment dose $s$ relative to a baseline (i.e., no treatment). Based on the proof for the binary treatment setting by Neal (2020), we prove that $\tau_{s}(\mathbf{x})$ is identifiable as (29) and estimable via causal machine learning methods under three assumptions: Consistency, Ignorability, Overlap (Section 4.1).
Proof.
From the definition of $\tau_{s}(\mathbf{x})$ and linearity of expectation:
$\tau_{s}(\mathbf{x})$ $=\mathbb{E}[Y(s)-Y(0)\mid\mathbf{X}=\mathbf{x}]$ (25)
$=\mathbb{E}[Y(s)\mid\mathbf{X}=\mathbf{x}]-\mathbb{E}[Y(0)\mid\mathbf{X}=\mathbf{x}]$ (26)
By Ignorability, potential outcomes are independent of treatment assignment given $\mathbf{X}$:
$\mathbb{E}[Y(s)\mid\mathbf{X}=\mathbf{x}]$ $=\mathbb{E}[Y(s)\mid\mathbf{X}=\mathbf{x},S=s].$ (27)
Applying Consistency to (27):
$\mathbb{E}[Y(s)\mid\mathbf{X}=\mathbf{x},S=s]$ $=\mathbb{E}[Y\mid\mathbf{X}=\mathbf{x},S=s].$ (28)
Combining (26)–(28):
$\tau_{s}(\mathbf{x})$ $=\mathbb{E}[Y\mid\mathbf{X}=\mathbf{x},S=s]-\mathbb{E}[Y\mid\mathbf{X}=\mathbf{x},S=0],$ (29)
where Overlap ensures the conditional expectations in (27)-(29) are well-defined. ∎
Appendix C Details regarding semi-synthetic data
The original IHDP dataset contains 747 observations and 25 features. Following Nie et al. (2021), synthetic counterfactuals are generated by Equations 30-33 where $S_{\mathrm{con}}=\bigl{\{}1,2,3,5,6\bigr{\}}$ is the index set of continuous features, $S_{\mathrm{bin},1}=\bigl{\{}4,7,8,9,10,11,12,13,14,15\bigr{\}}$ and $S_{\mathrm{bin},2}=\bigl{\{}16,17,18,19,20,21,22,23,24,25\bigr{\}}$ are two sets of binary features, $c_{1}=\mathbb{E}\left[\frac{\sum_{i\in S_{\mathrm{dis},1}}x_{i}}{|S_{\mathrm{bin},1}|}\right]$, and $c_{2}=\mathbb{E}\left[\frac{\sum_{i\in S_{\mathrm{dis},2}}x_{i}}{|S_{\mathrm{bin},2}|}\right]$.
@@ Appendix C Details regarding semi-synthetic data
Appendix D Number of dose bins $\delta$
@@
Appendix E Hyperparameter tuning
Method Hyperparameter Values #Models S-Learner (rf) Estimators {10, 20, 50, 100, 200, 500} 72 Criterion {Sq. error, Abs. error} Max depth {5, 15, None} Max features {Sqrt, Log2} S-Learner (mlp) Learning rate {0.001, 0.01} 128 L2 regularization {0.0, 0.1} Batch size {64, 128} Num layers {2, 3} Hidden size {32, 64} Steps {500, 1000, 2000, 5000} Optimizer {Adam} DRNet Learning rate {0.001, 0.01} 256 L2 regularization {0.0, 0.1} Batch size {64, 128} Num repr. layers {2, 3} Num inf. layers {1, 2} Hidden size {32, 64} Steps {500, 1000, 2000, 5000} Num dose strata {10} Optimizer {Adam} VCNet Learning rate {0.001, 0.01} 32 Batch size {64, 128} Hidden size {32, 64} Steps {500, 1000, 2000, 5000} Optimizer {Adam}
Table 5: Hyperparameter search space. The selected hyperparameters are underlined.
Appendix F More details on Experiment 1
F.1 Dose-response estimation
@@
F.2 Scalability
We provide additional runtime analysis for increasing dataset sizes. The dataset size is semi-synthetically scaled by a factor $N$ (with $N=1$ corresponding to the original dataset), by randomly oversampling observations and adding slight noise to ensure that each observation remains unique. We adopt the setup of Experiment 1 (Section 5.3.1), where the available budget is also scaled by $N$, and compare the runtime of the ILP solution with that of the heuristic approach. Figure 10 shows the runtime (in seconds) as a function of the problem size (measured by the scaling factor $N$). Both axes are shown on a logarithmic scale. The clear advantage of the heuristic method is that it scales well. However, it does not take into account any side constraints, which are allowed by the ILP. The downside of the ILP is that runtimes increase rapidly with the problem size, becoming impractical for large-scale instances. This highlights the classic trade-off: while the ILP allows for more expressive modeling and constraint handling, its computational cost means that problem size is limited. In contrast, the heuristic remains feasible in terms of computation time even as the dataset grows.
Figure 10: Runtime of the ILP and heuristic as a function of problem size. The original dataset size corresponds with $N=1$.